OpenStudy (anonymous):

Solve the equation. x - 9 = -6x + 5 A. 14 B. 7 C. 2 D. 21

OpenStudy (anonymous):

hmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmm

OpenStudy (anonymous):

wellz. You have to first find out what the x is. And how do you do that?

OpenStudy (anonymous):

I mean it's kinda easy but im makin u tell me ^-^

OpenStudy (anonymous):

Okay, this is simple. What is the first step to solving a two-step equation?

OpenStudy (anonymous):

x+6x= 5+9 7x=14 x=2

OpenStudy (anonymous):

Im horrible at this. =(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x-9=-6x+5 +6x +6x 7x-9=5 +9 +9 7 /7x=14/7 x=2